OpenStudy (anonymous):

A polynomial P(x) has rational coefficients. Name additional roots p(x) given the roots 5+sqrt 3 and - sqrt 2

OpenStudy (anonymous):

irrational roots of such a polynomial come in conjugate pairs so if \(a+\sqrt{b}\) is a root, so is \(a-\sqrt{b}\)

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

these are called "conjugates" the conjugate of \(a+\sqrt{b}\) is \(a-\sqrt{b}\) what, therefore, is the conjugate of \(5+\sqrt{3}\)?

OpenStudy (anonymous):

5-sqrt3

OpenStudy (anonymous):

yes, so that must be another root

OpenStudy (anonymous):

what about the conjugate of \(-\sqrt{2}\)? hint, it is the same as \(0-\sqrt{2}\)

OpenStudy (anonymous):

sqrt2

OpenStudy (anonymous):

yes, so you have the answer

OpenStudy (anonymous):

thanx so much!!!!!

OpenStudy (anonymous):

yw